# Multi-height analysis of rational points of toric varieties

### v1

@@ 2.5. The Möbius function associated to the universal torsor
For this section, we follow the notations of Salberger [18, Notations 11.8] in order to introduce the Möbius function associated to the universal torsor $\mathcal{T}$ over $X$.
@@
In [18, definition 11.9], Salberger defined a Möbius function:
@@
For a prime number $p$, we denote by $\sum\limits_{d}^{(p)}$ the sum over the $\Sigma(1)$-tuple of the form $d=(p^{e_{\rho}})_{\rho\in\Sigma(1)}$ where $(e_{\rho})_{\rho\in\Sigma(1)}$ belongs to $\mathbf{N}^{\Sigma(1)}$. We also define $\Pi(d)=\prod\limits_{\rho\in\Sigma(1)}d_{\rho}$ for any $d\in(\mathbf{N}^{*})^{\Sigma(1)}$. Salberger showed (see [18, 11.14]) the following result:
@@
Recall that we denote by $f$ be the smallest positive integer such that there exists $f$ rays of $\Sigma$ not contained in a cone of $\Sigma$. We have the following result from [18, lemma 11.15]:
@@
is absolutely convergent for $\Re(s)>\frac{1}{f}$. In the same way,
@@
for $\Re(s)>\frac{1}{f}$.
Using this proposition, one can deduce (see [18, lemma 11.19]) the following lemma:
@@
Let $\epsilon\in]0,1-\frac{1}{f}[$. We have the following upper bounds:
@@
    $\sum\limits_{\Pi(d)\leqslant Y}|\mu(d)|=O\left(Y^{\frac{1}{f}+\varepsilon}\right)$
@@
    $\sum\limits_{\Pi(d)\geqslant Y}\frac{|\mu(d)|}{\Pi(d)}=O\left(\frac{1}{Y^{1-\frac{1}{f}-\epsilon}}\right).$
@@ 3. Lift of the number of rational points of the toric variety to its universal torsor
The method of lifting the number of points to universal torsors have been used in a particular case in [13], then extended to toric varieties in [18, section 11] and generalised in [15],[17].
@@
$\sharp\Bigl{(}\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$
@@
In what follows, $\operatorname{Vol}$ means the usual euclidean volume given by the Lebesgue measures. The aim of this section is to relate $\sharp\Bigl{(}\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$ and the volume of $\mathcal{D}_{B}=\{y\in\mathcal{T}(\mathbf{R})\mid h_{\mathcal{T},\infty}(y)\in\mathrm{D}_{B}\}$. To estimate the difference between this two terms, we shall use a result of geometry of numbers due to Davenport [8]. The statement we shall use is from [12, proposition 24]:
@@
Each time we shall use this proposition in this article, the constant $m$, $k$ and $l$ will not vary. We shall apply this theorem to the semi-algebraic bounded set $\Phi(\mathcal{D}_{1})$ where:
@@
    $\Phi$ is the multiplication by an element $(a_{\rho})_{\rho\in\Sigma(1)}\in(\mathbf{R}_{+}^{\times})^{\Sigma(1)}$. In other words:
    $\Phi:$ $\mathbf{R}^{\Sigma(1)}\rightarrow\mathbf{R}^{\Sigma(1)}$
@@
Hence we shall have $m=\sharp\Sigma(1)$, $k$ depends only on the choice of $\mathrm{D}$, the finite union of compact polyhedrons and $l$ depends on the choice of $\mathrm{D}$ and on the maximum of the degree of the monomial $X^{D}$ for $D\in\mathcal{M}_{k}$ and $k\in\{1,..,t\}$ (see notations from 2.2). By this remark, applying the previous proposition gives that there exists $C_{1}>0$ and $C_{2}>0$, which does not depend on the choice of $\Phi$, such that:
@@
For the rest of the article, we shall suppose that $B\geqslant 1$.
3.1.1. A compacity result
@@
Theorem 3.2.
There exists a constant $C>0$ (depending on the compact subset $\mathrm{D}$) such that for any $y\in\mathcal{D}_{B}$, for any $\rho\in\Sigma(1)$, we have:
$|y_{\rho}|\leqslant CB^{\langle[D_{\rho}],u\rangle}.$
@@
$\mathcal{D}_{B}=B^{u}.\mathcal{D}_{1}.$
Then to prove the previous theorem, we have to recall that the action of $B^{u}$ on the coordinate associated to $\rho\in\Sigma(1)$ is given by multiplication by $B^{\langle[D_{\rho}],u\rangle}$ by proposition 2.17 and we can conclude using the following lemma:
@@
There exists a constant $C>0$ (depending on the compact subset $\mathrm{D}$) such that for any $y\in\mathcal{D}_{1}$, for any $\rho\in\Sigma(1)$, we have:
$|y_{\rho}|\leq C.$
@@
Recall that as we take $u\in(\text{C}_{\text{eff}}(X)^{\vee})^{\circ}$ i.e for any $L\in\text{C}_{\text{eff}}(X)$, $\langle L,u\rangle>0$, then we have that:
@@ 3.1.2. Upper bound for the number of points in the complementary of the torus
There exists a constant $C>0$ such that:
$\Big{|}\sharp\Bigl{(}\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\Big{|}\leqslant C\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@ 3.2. Extension of our estimation in order to apply the Möbius inversion
Denote by $\varphi:\mathbf{R}^{\Sigma(1)}\rightarrow\mathbf{R}^{\Sigma(1)}$ the diagonal endomorphism given by multiplication by $d_{\rho}$ the coordinate associated to the edge $\rho$. Recall that we have $\operatorname{covol}(\Lambda_{d})=\Pi(d)$. We can interpret $N_{d}(B)$ as follows:
$N_{d}(B)$ =$=\sharp\left\{y\in\mathbf{Z}^{\Sigma(1)}\mid h_{\mathcal{T},\infty}\left(\varphi(y)\right)\in\mathrm{D}_{B}\text{ and }\varphi(y)\in T_{\Sigma(1)}(\mathbf{R})\right\}$
=$=\sharp\left\{y\in\mathbf{Z}^{\Sigma(1)}\mid h_{\mathcal{T},\infty}\left(B^{-u}.\varphi(y)\right)\in\mathrm{D}\text{ and }B^{-u}.\varphi(y)\in T_{\Sigma(1)}(\mathbf{R})\right\}$
@@
$N_{d}(B)=\sharp\left(\mathbf{Z}^{\Sigma(1)}\cap\Phi(\mathcal{D}_{1})\right)$
where $\Phi(z)=\varphi^{-1}(B^{u}.z)$. Using this description, we begin by giving an upper bound of $N_{d}(B)$:
@@
Again by lemma 3.3, there exists $N>1$ such that $\mathcal{D}_{1}\subset\prod\limits_{\rho\in\Sigma(1)}\left([-N,N]-\{0\}\right)$. Hence $\Phi(\mathcal{D}_{1})\subset\prod\limits_{\rho\in\Sigma(1)}\left(\left[-N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}},N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}\right]-\{0\}\right)$. So we have that:
$N_{d}(B)=\sharp(\mathbf{Z}^{\Sigma(1)}\cap\Phi(\mathcal{D}_{1}))\leqslant 2^{\Sigma(1)}.\prod\limits_{\rho\in\Sigma(1)}\lfloor N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}\rfloor\leqslant(2.N)^{\sharp\Sigma(1)}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{\Pi(d)}.$
@@
There exists a constant $C>0$ (depending on the choice of $\mathrm{D})$ such that:
@@
Denote by $\Phi_{I}:\mathbf{R}^{I}\rightarrow\mathbf{R}^{I}$ the morphism which maps
$(x_{\rho})_{\rho\in I}\mapsto(\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}.x_{\rho})_{\rho\in I}.$
@@
$\operatorname{proj}_{I}\circ\Phi=\Phi_{I}\circ\operatorname{proj}_{I}.$
Hence we get, using this, that there exists $C>0$ (which depends only on the compact subset $\mathrm{D}$) such that for any non empty set $I\subsetneq\Sigma(1)$:
$\operatorname{Vol}(\operatorname{proj}_{I}(\Phi(\mathcal{D}))\leqslant C.\frac{B^{\langle\sum\limits_{\rho\in I}[D_{\rho}],u\rangle}}{\prod\limits_{\rho\in I}d_{\rho}}\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@ 3.3. The Möbius inversion
Firstly, by remark 3.11, the sum is finite. Now we have:
@@
Let $\varepsilon\in]0,1-\frac{1}{f}[$. There exists a constant $C>0$ such that:
@@
$\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{f}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@
Recall that by propositions 2.29 and 2.30, we have that:
@@
Let $\varepsilon\in]0,1-\frac{1}{f}[$. Then by lemma 2.31, we get that:
@@
$\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.B^{(\frac{1}{f}+\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}.$
@@
$\leqslant C^{\prime\prime}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{f}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@
Let $\varepsilon\in]0,1-\frac{1}{f}[$. There exists a constant $C>0$ such that:
$\left|(\mathcal{T}(\mathbf{Z}))_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p})).\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{f}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}$

### v2

@@ 3.5. The Möbius function associated to the universal torsor
For this section, we follow the notations of Salberger [20, Notations 11.8] in order to introduce the Möbius function associated to the universal torsor $\mathcal{T}$ over $X$.
@@
In [20, definition 11.9], Salberger defined a Möbius function:
@@
For a prime number $p$, we denote by $\sum\limits_{d}^{(p)}$ the sum over the $\Sigma(1)$-tuple of the form $d=(p^{e_{\rho}})_{\rho\in\Sigma(1)}$ where $(e_{\rho})_{\rho\in\Sigma(1)}$ belongs to $\mathbf{N}^{\Sigma(1)}$. We also define $\Pi(d)=\prod\limits_{\rho\in\Sigma(1)}d_{\rho}$ for any $d\in(\mathbf{N}^{*})^{\Sigma(1)}$. Salberger showed (see [20, 11.14]) the following result:
@@
Recall that we denote by $l$ be the smallest positive integer such that there exists $l$ rays of $\Sigma$ not contained in a cone of $\Sigma$. We have the following result from [20, lemma 11.15]:
@@
is absolutely convergent for $\Re(s)>\frac{1}{l}$. In the same way,
@@
for $\Re(s)>\frac{1}{l}$.
Using this proposition, one can deduce (see [20, lemma 11.19]) the following lemma:
@@
Let $\epsilon\in]0,1-\frac{1}{l}[$. We have the following upper bounds:
@@
    $\sum\limits_{\Pi(d)\leqslant Y}|\mu(d)|=O\left(Y^{\frac{1}{l}+\varepsilon}\right)$
@@
    $\sum\limits_{\Pi(d)\geqslant Y}\frac{|\mu(d)|}{\Pi(d)}=O\left(\frac{1}{Y^{1-\frac{1}{l}-\epsilon}}\right).$
@@ 4. Lift of the number of rational points of the toric variety to its universal torsor
The method of lifting the number of points to universal torsors have been used in a particular case in [15], then extended to toric varieties in [20, section 11] and generalised in [17],[19].
@@
$\sharp\Bigl(\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}.$
@@
In what follows, $\operatorname{Vol}$ means the usual euclidean volume given by the Lebesgue measures. The aim of this section is to relate $\sharp\Bigl(\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$ and the volume of $\mathcal{D}_{B}=\{y\in\mathcal{T}(\mathbf{R})\mid h_{\mathcal{T},\infty}(y)\in\mathrm{D}_{B}\}$. To estimate the difference between this two terms, we shall use a result of geometry of numbers due to Davenport [9]. The statement we shall use is from [14, proposition 24]:
@@
Each time we shall use this proposition in this article, the constant $m$, $k$ and $l$ will not vary. We shall apply this proposition to the semi-algebraic bounded set $\Psi(\mathcal{D}_{1})$ where:
@@
    $\Psi$ is the multiplication by an element $(a_{\rho})_{\rho\in\Sigma(1)}\in(\mathbf{R}_{+}^{\times})^{\Sigma(1)}$. In other words:
    $\Psi:\$$\mathbf{R}^{\Sigma(1)}\rightarrow\mathbf{R}^{\Sigma(1)}$
@@
Hence we shall have $m=\sharp\Sigma(1)$, $k$ depends only on the choice of $\mathrm{D}$, the finite union of compact polyhedrons and $l$ depends on the choice of $\mathrm{D}$ and on the maximum of the degree of the monomial $X^{D}$ for $D\in\mathcal{M}_{k}$ and $k\in\{1,..,t\}$ (see notations from 3.2). By this remark, applying the previous proposition gives that there exists $C_{1}>0$ and $C_{2}>0$, which does not depend on the choice of $\Phi$, such that:
@@
For the rest of the article, we shall suppose that $B\geqslant 1$. Recall we have chosen a vector $u\in\left(\text{C}_{\text{eff}}(X)^{\vee}\right)^{\circ}$ and defined $\mathrm{D}_{B}=\mathrm{D}+\log(B)u$.
4.1.1. An upper bound on the coordinates
@@
Proposition 4.2.
There exists a constant $C_{\mathrm{D}}>0$ (depending on the compact subset $\mathrm{D}$) such that for any $y\in\mathcal{D}_{B}$, for any $\rho\in\Sigma(1)$, we have:
$|y_{\rho}|\leqslant C_{\mathrm{D}}B^{\langle[D_{\rho}],u\rangle}.$
@@
$\mathcal{D}_{B}=B^{-u}.\mathcal{D}_{1}.$
Then to prove the previous proposition, we have to recall that the action of $B^{-u}$ on the coordinate associated to $\rho\in\Sigma(1)$ is given by multiplication by $B^{\langle[D_{\rho}],u\rangle}$ by proposition 3.19 and we can conclude using the following lemma:
@@
There exists a constant $C_{\mathrm{D}}>0$ (depending on the compact subset $\mathrm{D}$) such that for any $y\in\mathcal{D}_{1}$, for any $\rho\in\Sigma(1)$, we have:
$|y_{\rho}|\leq C_{\mathrm{D}}.$
@@
Recall that as we take $u\in(\text{C}_{\text{eff}}(X)^{\vee})^{\circ}$ i.e. for any $L\in\text{C}_{\text{eff}}(X)$, $\langle L,u\rangle>0$, then we have that:
@@ 4.1.2. Upper bound for the number of points in the complementary of the torus
There exists a constant $C_{\mathrm{D}}>0$ such that:
$\Big|\sharp\Bigl(\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\Big|\leqslant C_{\mathrm{D}}\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@ 4.2. Extension of our estimation in order to apply the Möbius inversion
Denote by $\psi:\mathbf{R}^{\Sigma(1)}\rightarrow\mathbf{R}^{\Sigma(1)}$ the diagonal endomorphism given by multiplication by $d_{\rho}$ the coordinate associated to the edge $\rho$. Recall that we have $\operatorname{covol}(\Lambda_{d})=\Pi(d)$. We can interpret $N_{d}(B)$ as follows:
$N_{d}(B)$ $=\sharp\left\{y\in\mathbf{Z}^{\Sigma(1)}\mid h_{\mathcal{T},\infty}\left(\psi(y)\right)\in\mathrm{D}_{B}\text{ and }\psi(y)\in T_{\Sigma(1)}(\mathbf{R})\right\}$
$=\sharp\left\{y\in\mathbf{Z}^{\Sigma(1)}\mid h_{\mathcal{T},\infty}\left(B^{u}.\psi(y)\right)\in\mathrm{D}\text{ and }B^{u}.\psi(y)\in T_{\Sigma(1)}(\mathbf{R})\right\}$
@@
$N_{d}(B)=\sharp\left(\mathbf{Z}^{\Sigma(1)}\cap\Psi(\mathcal{D}_{1})\right)$
where $\Psi(z)=\psi^{-1}(B^{-u}.z)$. Using this description, we begin by giving an upper bound of $N_{d}(B)$:
@@
Again by lemma 4.3, there exists $N>1$ such that $\mathcal{D}_{1}\subset\prod\limits_{\rho\in\Sigma(1)}\left([-N,N]-\{0\}\right)$. Hence $\Psi(\mathcal{D}_{1})\subset\prod\limits_{\rho\in\Sigma(1)}\left(\left[-N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}},N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}\right]-\{0\}\right)$. So we have that:
$N_{d}(B)=\sharp(\mathbf{Z}^{\Sigma(1)}\cap\Psi(\mathcal{D}_{1}))\leqslant 2^{\Sigma(1)}.\prod\limits_{\rho\in\Sigma(1)}\Big\lfloor N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}\Big\rfloor\leqslant(2.N)^{\sharp\Sigma(1)}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{\Pi(d)}.$
@@
There exists a constant $C_{\mathrm{D}}>0$ (depending on the choice of $\mathrm{D})$ such that:
$\left|N_{d}(B)-\frac{1}{\Pi(d)}.2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C_{\mathrm{D}}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@
Denote by $\Psi_{I}:\mathbf{R}^{I}\rightarrow\mathbf{R}^{I}$ the morphism which maps
$(x_{\rho})_{\rho\in I}\mapsto\left(\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}.x_{\rho}\right)_{\rho\in I}.$
@@
$\operatorname{proj}_{I}\circ\Psi=\Psi_{I}\circ\operatorname{proj}_{I}.$
Using this we get that there exists $C_{\mathrm{D}}>0$ (which depends only on the compact subset $\mathrm{D}$) such that for any non empty set $I\subsetneq\Sigma(1)$:
$\operatorname{Vol}(\operatorname{proj}_{I}(\Psi(\mathcal{D}))\leqslant C_{\mathrm{D}}.\frac{B^{\langle\sum\limits_{\rho\in I}[D_{\rho}],u\rangle}}{\prod\limits_{\rho\in I}d_{\rho}}\leqslant C_{\mathrm{D}}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@ 4.3. The Möbius inversion
Firstly, by remark 4.11, the sum is finite. Now using the definition of 3.30, we have:
@@
Let $\varepsilon\in]0,1-\frac{1}{l}[$. There exists a constant $C>0$ such that:
@@
$\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{l}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@
Recall that by propositions 3.31 and 3.32, we have that:
@@
Let $\varepsilon\in]0,1-\frac{1}{l}[$. Then by lemma 3.33, we get that:
@@
$\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.B^{(\frac{1}{l}+\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}.$
@@
$\leqslant C^{\prime\prime}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{l}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
@@
Let $\varepsilon\in]0,1-\frac{1}{l}[$. There exists a constant $C>0$ such that:
$\left|(\mathcal{T}(\mathbf{Z}))_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p})).\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{l}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}$